# Small-Time Local Controllability of the multi-input bilinear Schrödinger equation thanks to a quadratic term

### v3

@@ 4.1 Asymptotic estimates on the quadratic term of the solution
$B=\sum_{k=0}^{l}(-1)^{k}\left(\sum_{j=1}^{+\infty}\tilde{d_{j}}\nu_{j}^{k}\omega_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{j=1}^{+\infty}\tilde{d_{j}}\nu_{j}^{k}\omega_{j}^{k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1})\right)+\\ (-1)^{l+1}\sum_{j=1}^{+\infty}\tilde{d_{j}}\nu_{j}^{l+1}\omega_{j}^{l+1}I^{j}_{T}(v_{l+1},u_{l+1}).$ (4.9)
@@
Using the equations (3.12), (3.13), (3.14), (4.11), (4.12), and (4.13), we obtain (4.1). ∎
@@ 4.3 Motion along $\varphi_{K}$ and $i\varphi_{K}$
thanks to the remark 1.2 and the estimate (4.31). Using the equation (4.29) and (4.32), we obtain the desired inequality, because $\tau_{l}\leqslant 2\alpha_{n}+s_{n}(1-m-l)$. ∎
@@
Let $0<T_{1}<T$, one defines controls that allow to move along $i^{n}\varphi_{K}$, and correcting the linear part. More precisely, for $z\in\mathbb{R}$,
@@
where $\mathcal{H}$ is defined in (1.12). Thus, we just have to estimate $|\left\langle\psi(T;(U_{z},V_{z}),\varphi_{1}),\psi_{K}(T)\right\rangle-i^{n}z|$. Using the triangular inequality,
@@
We choose $0<\varepsilon<\frac{1}{48}$. We estimate these terms
@@
    $\begin{split}\left\|(U_{z},V_{z})\right\|_{L^{1}}^{3}&\leqslant C|z|^{3\alpha_{n}+\frac{3}{2}s_{n}}+C|z|^{3\tau_{0}}\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}+\left(\frac{1}{16}-3\varepsilon s_{n}\right)+\frac{9}{4}s_{n}}\\ &\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}+\left(\frac{1}{16}-3\varepsilon\right)}\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}}\leqslant C|z|^{1+\frac{1}{5}s_{n}}.\end{split}$
@@
We consider $(u_{z},v_{z})_{z\in\mathbb{R}}$ the family of controls associated with $i^{n}\varphi_{K}$ in Proposition 4.8. First, we show that there exists $C>0$, for all $(\alpha,\beta)\in\mathbb{R}^{2}$, small enough,
@@
In the interval $[T,2T]$, the evolution of the system is free ($(u,v)\equiv 0$), the solution is given by $\psi(2T)=e^{-iA(2T-T)}\psi(T)$. Using that $e^{-iAT}$ is an isometry from $H^{2(p+m)+3}_{(0)}$ to $H^{2(p+m)+3}_{(0)}$,
@@
Finally, as $e^{-iAT}(i^{n}\alpha\psi_{K}(2T))=i^{n}\alpha\psi_{K}(3T)$, using (4.39) with $\beta$ instead of $\alpha$, we obtain (4.38). Thus, for $T\in\left(0,\frac{\pi}{2\omega_{K}}\right)$, and $z\in\mathbb{R}$, taking $\beta=\frac{z}{\sin(2\omega_{K}T)}$ and $\alpha=-\beta\cos(2\omega_{k}T)$, we conclude. ∎
@@ 4.4 Conclusion
We use Theorem 1.7. More precisely, we consider $X=H^{2(p+m)+3}_{(0)}(0,1)$ and $E_{T}=H^{m}_{0}(0,T)^{2}$. Moreover, for all $T>0$,
@@ 5. $(A_{5})$
    Finally, as $\langle\mu_{1}\varphi_{1},\varphi_{1}\rangle\neq 0$, Propositions 4.8 and 4.9 show that $\varphi_{K}$ and $i\varphi_{K}$ are small-time continuously approximately reachable vectors. Then, as a supplementary space of $H$ is given by $\text{Span}_{\mathbb{R}}(\varphi_{K},i\varphi_{K})$, this condition is verified.
@@ A.1 Existence of $\mu_{1},\mu_{2}$ verifying the hypotheses
This section is inspired by [18] and [16].
@@
    $\gamma_{2n+1}=-\langle[\underline{\operatorname{ad}}_{A}^{n}(\mu_{2}),\underline{\operatorname{ad}}_{A}^{n+1}(\mu_{1})]\varphi_{1},\varphi_{K}\rangle=2(-1)^{n+1}\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle+Q_{n}(\mu_{1},\mu_{2}),$
@@
    $\gamma_{2n}=\langle[\underline{\operatorname{ad}}_{A}^{n}(\mu_{1}),\underline{\operatorname{ad}}_{A}^{n}(\mu_{2})]\varphi_{1},\varphi_{K}\rangle=4n(-1)^{n}\langle\mu_{1}^{(4n-2)}\mu^{\prime}_{2}\varphi_{1}^{\prime},\varphi_{K}\rangle+Q_{n}(\mu_{1},\mu_{2}),$
@@
Thus, thanks (A.2), one obtains
@@
$\left|\langle\mu_{1}^{(2(n+1)+i)}\mu_{2}^{(2n-i)}\varphi_{1},\varphi_{K}\rangle-(-1)^{i}\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n+1}}\left\|\mu_{2}\right\|_{L^{2}},$ (A.4)
$\left|\langle\mu_{1}^{(2(n+1)-i)}\mu_{2}^{(2n+i)}\varphi_{1},\varphi_{K}\rangle-(-1)^{i}\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n+1}}\left\|\mu_{2}\right\|_{L^{2}}.$ (A.5)
Using (A.3), (A.4) and (A.5) one has
$\left|\gamma_{2n+1}+\left(\alpha_{0}^{n+1}\sum_{i=0}^{n}\alpha_{i}^{n}(-1)^{i}-\alpha_{0}^{n}\sum_{i=0}^{n+1}\alpha_{i}^{n+1}(-1)^{i}\right)\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle\right|\\ \leqslant\left\|\mu_{1}\right\|_{H^{4n+1}}\left\|\mu_{2}\right\|_{L^{2}}.$
@@
$\left|\gamma_{2n}-\alpha_{0}^{n}\displaystyle\sum_{i=0}^{n}\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i)}\right)\varphi_{1},\varphi_{K}\right\rangle-\right.\\ \left.\alpha_{0}^{n}\sum_{i=1}^{n}i\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i-1)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i-1)}\right)\varphi_{1}^{\prime},\varphi_{K}\right\rangle-\right.\\ \left.\alpha_{1}^{n}\sum_{i=0}^{n}\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i-1)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i-1)}\right)\varphi_{1}^{\prime},\varphi_{K}\right\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n-3}}\left\|\mu_{2}\right\|_{H^{1}}.$$$
Thus, we obtain with same estimates
$\left|\gamma_{2n}-\left(2\alpha_{1}^{n}\sum_{i=0}^{n}\alpha_{i}^{n}(-1)^{i}\right)\langle\mu_{1}^{(4n-2)}\mu_{2}^{\prime}\varphi_{1}^{\prime},\varphi_{K}\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n-3}}\left\|\mu_{2}\right\|_{H^{1}}.$
By induction, we obtain $\alpha_{1}^{n}=2n.$ This proves the result. ∎
@@
Let $K\in\mathbb{N}$, $K\geqslant 2$ and $\bar{x}\in(0,1)$ such that $\sin(K\pi\bar{x})=0$. There exists $\delta>0$ such that for every $n\in\mathbb{N}^{*}$, for every $J^{+}$ (resp. $J^{-}$) open interval of $(\bar{x},\bar{x}+\delta)$ (resp. $(\bar{x}-\delta,\bar{x})$), there exist $\mu_{1,+}^{n},\mu_{2,+}^{n}$ (resp. $\mu_{1,-}^{n},\mu_{2,-}^{n}$) in $\mathcal{C}_{c}^{\infty}(0,1)$ supported on $J^{+}$ (resp. $J^{-}$) such that:
$\langle\mu_{1,\pm}^{n}\varphi_{1},\varphi_{K}\rangle=\langle\mu_{2,\pm}^{n}\varphi_{1},\varphi_{K}\rangle=\gamma_{1}(\mu_{1,\pm}^{n},\mu_{2,\pm}^{n})=\cdots=\gamma_{n-1}(\mu_{1,\pm}^{n},\mu_{2,\pm}^{n})=0,$
@@
$\gamma_{n}(\mu_{1,\pm}^{n},\mu_{2,\pm}^{n})=\pm 1.$
@@
Let $n\geqslant 1$, $K\geqslant 2$, $p,m\geqslant 0$, such that $\lfloor\frac{n}{2}\rfloor\leqslant p$. There exist $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{lin,K,1}}$, $\mathbf{(H)_{lin,K,2}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ and $\mathbf{(H)_{quad,K,4}}$.
@@ A.2.1 Definition of Lie brackets in infinite dimension
$\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi=\sum_{l=0}^{k}\alpha_{l}^{k}\mu^{(2k-l)}\varphi^{(l)},$
@@
    We now prove that $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi\in H^{2}_{(0)}(0,1)$. The induction hypothesis gives: $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi=\displaystyle\sum_{l=0}^{k}\alpha_{l}^{k}\mu_{1}^{(2k-l)}\varphi^{(l)}.$ Then, one has $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\in H^{2}(0,1)$. Moreover, for all $l\in\llbracket 0,k\rrbracket$,
@@
    Then, the bracket is well-defined. The proposition is true by induction on $k$.
@@
    $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi\in H^{2k}_{(0)}(0,1)$. One recalls: $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi=\displaystyle\sum_{l=0}^{k-1}\alpha_{l}^{k-1}\mu_{1}^{(2k-l-2)}\varphi^{(l)}$. According to the regularity of $\varphi$ and $\mu_{1}$, this expansion gives $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi\in H^{2k}(0,1)$. Then, using Leibniz formula, for all $j\in\llbracket 0,k-1\rrbracket$, $l\in\llbracket 0,k-1\rrbracket$,
@@
Consequently, the operator $\left[\underline{\operatorname{ad}}_{A}^{k-1}(B_{1}),\underline{\operatorname{ad}}_{A}^{k}(B_{1})\right]$ is well-defined. ∎
Then, if $12\lfloor\frac{n+1}{2}\rfloor-3\leqslant p-$ 1, Proposition A.6 ensures that the brackets in (1.7) and (1.8) are well-defined.

### v4

@@ 4.1 Asymptotic estimates on the quadratic term of the solution
$B=\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k+1}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k+1})+\\ (-1)^{l+1}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(v_{l+1},u_{l+1}).$ (4.9)
@@
With the two Lemmas 4.2 and 4.4, we are now able to prove the main result of this subsection.
@@
Using the equations (3.14), (4.11), (4.12), and (4.13), we obtain (4.1). ∎
@@ 4.3 Motion along $\varphi_{K}$ and $i\varphi_{K}$
thanks to Remark 1.2 and the estimate (4.31). Using the equation (4.29) and (4.32), we obtain the desired inequality, because $\tau_{l}\leqslant 2\alpha_{n}+s_{n}(1-m-l)$. ∎
@@
Let $0<T_{1}<T$, one defines controls that allow to move along $i^{n}\varphi_{K}$, and correcting the linear part, thanks to hypothesis $\mathbf{(H)_{lin,K,2}}$. More precisely, for $z\in\mathbb{R}$,
@@
where $\mathcal{H}$ is defined in (1.13). Thus, we just estimate $|\left\langle\psi(T;(U_{z},V_{z}),\varphi_{1}),\psi_{K}(T)\right\rangle-i^{n}z|$. Using the triangular inequality,
@@
We choose $0<\varepsilon<\frac{1}{48}$. We estimate these terms.
@@
    $\begin{split}\left\|(U_{z},V_{z})\right\|_{L^{2}}^{3}&\leqslant C|z|^{3\alpha_{n}+\frac{3}{2}s_{n}}+C|z|^{3\tau_{0}}\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}+\left(\frac{1}{16}-3\varepsilon s_{n}\right)+\frac{9}{4}s_{n}}\\ &\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}+\left(\frac{1}{16}-3\varepsilon\right)}\leqslant C|z|^{1+\frac{1}{5}s_{n}}+C|z|^{\frac{17}{16}}\leqslant C|z|^{1+\frac{1}{5}s_{n}}.\end{split}$
@@
We consider $(u_{z},v_{z})_{z\in\mathbb{R}}$ the family of controls associated with $i^{n}\varphi_{K}$ in Proposition 4.8. First, we show that there exists $C>0$ such that for all $(\alpha,\beta)\in\mathbb{R}^{2}$, small enough,
@@
In the interval $[T,2T]$, the evolution of the system is free ($(u,v)\equiv 0$). Consequently, the solution is given by $\psi(2T)=e^{-iA(2T-T)}\psi(T)$. Using that $e^{-iAT}$ is an isometry from $H^{2(p+m)+3}_{(0)}$ to $H^{2(p+m)+3}_{(0)}$, we obtain
@@
Finally, as $e^{-iAT}(i^{n}\alpha\psi_{K}(2T))=i^{n}\alpha\psi_{K}(3T)$, using (4.39) with $\beta$ instead of $\alpha$, we obtain (4.38). Thus, for $T\in\left(0,\frac{\pi}{2\omega_{K}}\right)$, and $z\in(-\rho,\rho)$, taking $\beta=\frac{z}{\sin(2\omega_{K}T)}$ and $\alpha=-\beta\cos(2\omega_{k}T)$, we conclude. ∎
@@ 4.4 Conclusion
We use Theorem 1.7. More precisely, we define $X=H^{2(p+m)+3}_{(0)}(0,1)$ and $E_{T}=H^{m}_{0}(0,T)^{2}$. Moreover, for all $T>0$,
@@
We have to check all the assumptions of Theorem 1.7 to obtain the controllability result.
@@ 5. $(A_{5})$
    Finally, as $\langle\mu_{1}\varphi_{1},\varphi_{1}\rangle\neq 0$, Propositions 4.8 and 4.9 show that $\varphi_{K}$ and $i\varphi_{K}$ are small-time $H_{0}^{m}$-continuously approximately reachable vectors. Then, as a supplementary space of $H$ is given by $\text{Span}_{\mathbb{R}}(\varphi_{K},i\varphi_{K})$, this condition is verified.
@@ A.1 Existence of $\mu_{1},\mu_{2}$ verifying the hypotheses
This section is inspired by [16, 18]. We recall that the operator $A$ is defined in (1.2).
@@
    $\gamma_{2n+1}=-\langle[\underline{\operatorname{ad}}_{A}^{n+1}(\mu_{1}),\underline{\operatorname{ad}}_{A}^{n}(\mu_{2})]\varphi_{1},\varphi_{K}\rangle=2(-1)^{n}\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle+Q_{n}(\mu_{1},\mu_{2}),$
@@
    $\gamma_{2n}=\langle[\underline{\operatorname{ad}}_{A}^{n}(\mu_{1}),\underline{\operatorname{ad}}_{A}^{n}(\mu_{2})]\varphi_{1},\varphi_{K}\rangle=4n(-1)^{n}\langle\mu_{1}^{(4n-2)}\mu^{\prime}_{2},\varphi_{1}^{\prime}\varphi_{K}-\varphi_{K}^{\prime}\varphi_{1}\rangle+Q_{n}(\mu_{1},\mu_{2}),$
@@
Thus, thanks to (A.2), one obtains
@@
$\left|\langle\mu_{1}^{(2(n+1)\pm i)}\mu_{2}^{(2n\mp i)}\varphi_{1},\varphi_{K}\rangle-(-1)^{i}\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n+1}}\left\|\mu_{2}\right\|_{L^{2}}.$ (A.4)
Using (A.3) and (A.4), one has
$\left|\gamma_{2n+1}-\left(\alpha_{0}^{n+1}\sum_{i=0}^{n}\alpha_{i}^{n}(-1)^{i}-\alpha_{0}^{n}\sum_{i=0}^{n+1}\alpha_{i}^{n+1}(-1)^{i}\right)\langle\mu_{1}^{(4n+2)}\mu_{2}\varphi_{1},\varphi_{K}\rangle\right|\\ \leqslant\left\|\mu_{1}\right\|_{H^{4n+1}}\left\|\mu_{2}\right\|_{L^{2}}.$
@@
$\left|\gamma_{2n}-\alpha_{0}^{n}\displaystyle\sum_{i=0}^{n}\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i)}\right)\varphi_{1},\varphi_{K}\right\rangle-\right.\\ \left.\alpha_{0}^{n}\sum_{i=1}^{n}i\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i-1)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i-1)}\right),\varphi_{1}^{\prime}\varphi_{K}\right\rangle-\right.\\ \left.\alpha_{1}^{n}\sum_{i=0}^{n}\alpha_{i}^{n}\left\langle\left(\mu_{1}^{(2n-i)}\mu_{2}^{(2n+i-1)}-\mu_{2}^{(2n-i)}\mu_{1}^{(2n+i-1)}\right),\varphi_{1}^{\prime}\varphi_{K}\right\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n-3}}\left\|\mu_{2}\right\|_{H^{1}}.$$$
Hence, using similar estimates, we obtain the simpler expression
$\left|\gamma_{2n}+2\alpha_{0}^{n}\left(\sum_{i=0}^{n}i\alpha_{i}^{n}(-1)^{i}\right)\langle\mu_{1}^{(4n-2)}\mu_{2}^{\prime},(\varphi_{1}\varphi_{K})^{\prime}\rangle-\right.\\ \left.\left(2\alpha_{0}^{n}\sum_{i=0}^{n}i\alpha_{i}^{n}(-1)^{i}+2\alpha_{1}^{n}\sum_{i=0}^{n}\alpha_{i}^{n}(-1)^{i}\right)\langle\mu_{1}^{(4n-2)}\mu_{2}^{\prime},\varphi_{1}^{\prime}\varphi_{K}\rangle\right|\leqslant\left\|\mu_{1}\right\|_{H^{4n-3}}\left\|\mu_{2}\right\|_{H^{1}}.$
By induction, we obtain $\alpha_{1}^{n}=2n.$ The sums calculated in Lemma A.1 lead to the result. ∎
@@
Let $K\in\mathbb{N}$, $K\geqslant 2$. There exist $\mu_{1}^{n},\mu_{2}^{n}\in\mathcal{C}_{c}^{\infty}(0,1)$ such that
$\langle\mu_{1}^{n}\varphi_{1},\varphi_{K}\rangle=\langle\mu_{2}^{n}\varphi_{1},\varphi_{K}\rangle=\gamma_{1}(\mu_{1}^{n},\mu_{2}^{n})=\cdots=\gamma_{n-1}(\mu_{1}^{n},\mu_{2}^{n})=0,$
@@
$\gamma_{n}(\mu_{1}^{n},\mu_{2}^{n})=1.$
Remark A.4.
Choosing $\tilde{\mu_{1}^{n}}=-\mu_{1}^{n}$, one has $\gamma_{n}(\tilde{\mu_{1}^{n}},\mu_{2}^{n})=-1$.
@@
Let $n\geqslant 1$, $K\geqslant 2$, $p,m\geqslant 0$, such that $n\leqslant p+1$. There exist $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{lin,K,1}}$, $\mathbf{(H)_{lin,K,2}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ and $\mathbf{(H)_{quad,K,4}}$.
@@ A.2.1 Definition of Lie brackets in infinite dimension
$\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi=\sum_{l=0}^{k}\alpha_{l}^{k}\mu_{1}^{(2k-l)}\varphi^{(l)},$
@@
    We now prove that $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi\in H^{2}_{(0)}(0,1)$. By induction hypothesis, one has the equality $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi=\displaystyle\sum_{l=0}^{k}\alpha_{l}^{k}\mu_{1}^{(2k-l)}\varphi^{(l)}.$ Then, one has $\underline{\operatorname{ad}}_{A}^{k}(B_{1})\varphi\in H^{2}(0,1)$. Moreover, for all $l\in\llbracket 0,k\rrbracket$,
@@
    Then, the bracket is well-defined in $L^{2}(0,1)$. By induction, we obtain the result.
@@
    $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi\in H^{2k}_{(0)}(0,1)$. One recalls that $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi=\displaystyle\sum_{l=0}^{k-1}\alpha_{l}^{k-1}\mu_{1}^{(2k-l-2)}\varphi^{(l)}$. According to the regularity of $\varphi$ and $\mu_{1}$, this expansion gives $\underline{\operatorname{ad}}_{A}^{k-1}(B_{1})\varphi\in H^{2k}(0,1)$. Then, using Leibniz formula, for all $j\in\llbracket 0,k-1\rrbracket$, $l\in\llbracket 0,k-1\rrbracket$,
@@
Consequently, the operator $\left[\underline{\operatorname{ad}}_{A}^{k-1}(B_{1}),\underline{\operatorname{ad}}_{A}^{k}(B_{1})\right]$ is well-defined in $L^{2}(0,1)$. ∎
Then, if $12\lfloor\frac{n+1}{2}\rfloor-3\leqslant p-$ 1, Proposition A.7 ensures that the brackets in (1.8) and (1.9) are well-defined in $L^{2}(0,1)$.